OpenStudy (anonymous):

I WILL GIVE MEDALS Natalia constructed a right triangle with two side lengths measuring 3 and 5 units. She constructed a second right triangle similar to the first one, with the two corresponding sides measuring 5.1 and x units in length, respectively. What is the missing length, x, of the second triangle?

Mehek (mehek14):

write proportions \(\dfrac{5.1}{3} = \dfrac{x}{5}\)

Mehek (mehek14):

now cross multiply 5.1 x 5 = ?

Mehek (mehek14):

@DenishaH ?

Mehek (mehek14):

can you multiply 5.1 x 5?

OpenStudy (anonymous):

Yeah sorry I was gone xD. I got 25.5

Mehek (mehek14):

ok so what is \(\dfrac{25.5}{3}\)

OpenStudy (anonymous):

8.5

Mehek (mehek14):

so x = 8.5

OpenStudy (anonymous):

Thank you so much

Mehek (mehek14):

yw